whats the slope of
2x-6y=18

Answer by Alan3354(69443)

whats the slope of
2x-6y=18
--------------
Put the eqn in slope-intercept form.
That means solve for y.
2x-6y=18
6y = 2x - 18
y = (1/3)x - 3
---
The slope is the coefficient of x.
m = 1/3